# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ Appendix A Objective Function via Denoising Score Matching
Appendix B Proofs of the Results for the Multivariate Gaussian Initial Data with Unknown Mean
@@
In this section, we provide the proof of Theorem 3.1. We start by introducing the results which will be used in the proof of Theorem 3.1.
B.1 Preliminary Estimates
@@
We provide the results for the SGLD algorithm (15) with $\lambda$ given in (17), $\beta>0$, $C_{\mathsf{SGLD},1}$ and $C_{\mathsf{SGLD},2}$ given in Table 1, as well as for the auxiliary process (9), the discrete process (10), and the continuous-time interpolation (11) with $s$ given in (13) and $\gamma\in(0,1/2]$.
Proposition B.1.
@@
Proposition B.1 is derived from the definition of the stochastic gradient (14).
The proof of the following lemmas are postponed to Section B.3.
Lemma B.2.
@@
As a consequence of Lemma B.2, we derive the following corollary.
Corollary B.3.
@@
B.2 Proof of the Main Result in the Motivating Example
Proof of Theorem 3.1.
We derive the non-asymptotic estimate for $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ using the splitting
$\begin{split}W_{2}(\mathscr{L}(Y_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})&\leq W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K+1}}))+W_{2}(\mathscr{L}(Y_{t_{K+1}}),\mathscr{L}(\widetilde{Y}_{t_{K+1}}))\\ &\quad+W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K+1}}),\mathscr{L}(Y_{t_{K+1}}^{\text{aux}}))+W_{2}(\mathscr{L}(Y_{t_{K+1}}^{\text{aux}}),\mathscr{L}(Y_{K+1}^{\text{EM}})).\end{split}$ (29)
Since $t_{K+1}=T$, we have $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K+1}}))=0$. We provide upper bounds on the error made by approximating the initial condition of the backward process $Y_{0}\sim\mathscr{L}(X_{T})$ with $\widetilde{Y}_{0}\sim\pi_{\infty}$, i.e. $W_{2}(\mathscr{L}(Y_{t_{K+1}}),\mathscr{L}(\widetilde{Y}_{t_{K+1}}))$, the error made by approximating the score function with $s$, i.e. $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K+1}}),\mathscr{L}(Y_{t_{K+1}}^{\text{aux}}))$, and the discretisation error, i.e. $W_{2}(\mathscr{L}(Y_{t_{K+1}}^{\text{aux}}),\mathscr{L}(Y_{K+1}^{\text{EM}}))$, separately.
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K+1}}),\mathscr{L}(\widetilde{Y}_{t_{K+1}})).$
Applying Itô’s formula and using (3) and (4) with the score function given in (12), we have, for any $t\in[0,T]$,
@@ B.2 Proof of the Main Result in the Motivating Example
Integrating and taking expectation both sides in (30) yields
@@
Using (32), we have
$\begin{split}W_{2}(\mathscr{L}(Y_{t_{K+1}}),\mathscr{L}(\widetilde{Y}_{t_{K+1}}))&\leq\sqrt{\mathbb{E}[|Y_{t_{K+1}}-\widetilde{Y}_{t_{K+1}}|^{2}]}\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d}).\end{split}$ (33)
Upper bound on $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K+1}}),\mathscr{L}(Y_{t_{K+1}}^{\text{aux}})).$
Applying Itô’s formula and using the process (4) with the score function (12) and the process (9) with the approximating function (13), we have, for any $t\in[0,T]$,
@@
Integrating and taking expectation on both sides in (34) and using that the minimiser $\theta^{\star}=\mu$, we have
@@
Differentiating both sides of (35) and using Young’s inequality, we obtain
@@
Integrating both sides and using Lemma B.2 yields
@@
Using (36), we have
$W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K+1}}),\mathscr{L}(Y_{t_{K+1}}^{\text{aux}}))$ $\leq\sqrt{\mathbb{E}[|\widetilde{Y}_{t_{K+1}}-Y_{t_{K+1}}^{\text{aux}}|^{2}]}$ (37)
$\leq\sqrt{4/3}(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2})^{1/2}.$
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K+1}}^{\text{aux}}),\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}})).$
Applying Itô’s formula and using the processes (9) and (11) with the approximating function $s$ given in (13), we have, for any $t\in[0,T]$,
@@
Integrating both sides and taking expectation in (38) yields
@@
Differentiating both sides in (39), using Young’s inequality and $m_{T-t}-m_{T-\lfloor t/\gamma\rfloor\gamma}\leq\gamma m_{T-t}$, we have
@@
We derive an upper bound for the second term on the right-hand side in (40). Using Cauchy-Schwarz inequality, Itô formula applied to $tB_{t}$ and Young’s inequality, we have
$-2\mathbb{E}\left[\left\langle Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}},\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$=-2\mathbb{E}\left[\left\langle(Y_{t}^{\text{aux}}-Y_{\lfloor t/\gamma\rfloor\gamma}^{\text{aux}})-(\widehat{Y}_{t}^{\text{EM}}-\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}),\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$=-2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(-Y_{s}^{\text{aux}}+2m_{T-s}\hat{\theta})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$\quad+2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(-\widehat{Y}_{\lfloor s/\gamma\rfloor\gamma}^{\text{EM}}+2m_{T-\lfloor s/\gamma\rfloor\gamma}\hat{\theta})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$=2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(Y_{s}^{\text{aux}}-\widehat{Y}_{\lfloor s/\gamma\rfloor\gamma}^{\text{EM}})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$=2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(Y_{s}^{\text{aux}}-Y_{\lfloor s/\gamma\rfloor\gamma}^{\text{aux}})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$=2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(-Y_{\nu}^{\text{aux}})\mathrm{d}\nu\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$\quad+2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\sqrt{2}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\mathrm{d}\overline{B}_{\nu}\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$\leq 2\sqrt{2}\left(\mathbb{E}\left[\left|\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(-Y_{\nu}^{\text{aux}})\mathrm{d}\nu\mathrm{d}s\right|^{2}\right]\right)^{1/2}\left(\mathbb{E}\left[\left|\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right|^{2}\right]\right)^{1/2}$
$+4\mathbb{E}\left[\left\langle t\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}-\int_{\lfloor t/\gamma\rfloor\gamma}^{t}s\mathrm{d}\overline{B}_{s},\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]$
$\leq\sqrt{2}\gamma^{5/2}\left(\sup_{t\geq 0}\mathbb{E}\left[|Y_{t}^{\text{aux}}|^{2}\right]+d\right)+2d\gamma^{2}.$ (41)
Substituting (B.2) into (40) yields
@@
Integrating both sides in (42) and using Lemma B.4, Lemma B.5 and Corollary B.3, yield
@@
By using (43), we have
$W_{2}(\mathscr{L}(Y_{t_{K+1}}^{\text{aux}}),\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}))$ (44)
@@
Final upper bound on $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}}).$
Substituting (33), (37), (44) into (29) yields
$\begin{split}&W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d})\\ &\quad+\sqrt{4/3}(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2})^{1/2}\\ &\quad+\gamma(18d+132(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2}+|\theta^{*}|^{2}))^{1/2}\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d})\\ &\quad+(\sqrt{4/3}+2\sqrt{33})(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2})^{1/2}\\ &\quad+\gamma(18d+132|\theta^{*}|^{2})^{1/2}.\end{split}$ (45)
The bound for $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ in (45) can be made arbitrarily small by appropriately choosing parameters including $T,\beta,\lambda,n$ and $\gamma$. More precisely, for any $\delta>0$, we first choose $T>T_{\delta}$ with $T_{\delta}$ given explicitly in Table 1 such that
@@
Using (46), (47) and (48) in (45), we obtain $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})<\delta$. ∎
B.3 Proof of the Preliminary Results
We provide the proofs of the results of Appendix B.1.
Proof of Lemma B.2.
By (15), we have for any $n\in\mathbb{N}_{0}$,

### v5

@@ Appendix B Additional Discussions about Assumption 4
Appendix C Proofs of the Results for the Multivariate Gaussian Initial Data with Unknown Mean
@@
In this section, we provide the proof of Theorem 1. We start by introducing the results which will be used in the proof of Theorem 1.
C.1 Preliminary Estimates
@@
We provide the results for the SGLD algorithm 15 with $\lambda$ given in 17, $\beta>0$, $C_{\mathsf{SGLD},1}$ and $C_{\mathsf{SGLD},2}$ given in Table 1, as well as for the auxiliary process 9, the discrete process 10, and the continuous-time interpolation 11 with $s$ given in 13 and $\gamma\in(0,1/2]$.
Proposition 13.
@@
Proposition 13 is derived from the definition of the stochastic gradient 14.
The proof of the following lemmas are postponed to Section C.3.
Lemma 14.
@@
As a consequence of Lemma 14, we derive the following corollary.
Corollary 15.
@@
C.2 Proof of the Main Result in the Motivating Example
Proof of Theorem 1.
We derive the non-asymptotic estimate for $W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ using the splitting
$\begin{split}W_{2}(\mathcal{L}(Y_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})&\leq W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{K+1}}))+W_{2}(\mathcal{L}(Y_{t_{K+1}}),\mathcal{L}(\widetilde{Y}_{t_{K+1}}))\\ &\quad+W_{2}(\mathcal{L}(\widetilde{Y}_{t_{K+1}}),\mathcal{L}(Y_{t_{K+1}}^{\text{aux}}))+W_{2}(\mathcal{L}(Y_{t_{K+1}}^{\text{aux}}),\mathcal{L}(Y_{K+1}^{\text{EM}})).\end{split}$ (34)
Since $t_{K+1}=T$, we have $W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{K+1}}))=0$. We provide upper bounds on the error made by approximating the initial condition of the backward process $Y_{0}\sim\mathcal{L}(X_{T})$ with $\widetilde{Y}_{0}\sim\pi_{\infty}$, i.e. $W_{2}(\mathcal{L}(Y_{t_{K+1}}),\mathcal{L}(\widetilde{Y}_{t_{K+1}}))$, the error made by approximating the score function with $s$, i.e. $W_{2}(\mathcal{L}(\widetilde{Y}_{t_{K+1}}),\mathcal{L}(Y_{t_{K+1}}^{\text{aux}}))$, and the discretisation error, i.e. $W_{2}(\mathcal{L}(Y_{t_{K+1}}^{\text{aux}}),\mathcal{L}(Y_{K+1}^{\text{EM}}))$, separately.
Upper bound on $W_{2}(\mathcal{L}(Y_{t_{K+1}}),\mathcal{L}(\widetilde{Y}_{t_{K+1}})).$
Applying Itô’s formula and using 3 and 4 with the score function given in 12, we have, for any $t\in[0,T]$,
@@ C.2 Proof of the Main Result in the Motivating Example
Integrating and taking expectation both sides in 35 yields
@@
Using 37, we have
$\begin{split}W_{2}(\mathcal{L}(Y_{t_{K+1}}),\mathcal{L}(\widetilde{Y}_{t_{K+1}}))&\leq\sqrt{\mathbb{E}[|Y_{t_{K+1}}-\widetilde{Y}_{t_{K+1}}|^{2}]}\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d}).\end{split}$ (38)
Upper bound on $W_{2}(\mathcal{L}(\widetilde{Y}_{t_{K+1}}),\mathcal{L}(Y_{t_{K+1}}^{\text{aux}})).$
Applying Itô’s formula and using the process 4 with the score function 12 and the process 9 with the approximating function 13, we have, for any $t\in[0,T]$,
@@
Integrating and taking expectation on both sides in 39 and using that the minimiser $\theta^{\star}=\mu$, we have
@@
Differentiating both sides of 40 and using Young’s inequality, we obtain
@@
Integrating both sides and using Lemma 14 yields
@@
Using 41, we have
$\begin{split}W_{2}(\mathcal{L}(\widetilde{Y}_{t_{K+1}}),\mathcal{L}(Y_{t_{K+1}}^{\text{aux}}))&\leq\sqrt{\mathbb{E}[|\widetilde{Y}_{t_{K+1}}-Y_{t_{K+1}}^{\text{aux}}|^{2}]}\\ &\leq\sqrt{4/3}(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2})^{1/2}.\end{split}$ (42)
Upper bound on $W_{2}(\mathcal{L}(Y_{t_{K+1}}^{\text{aux}}),\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}})).$
Applying Itô’s formula and using the processes 9 and 11 with the approximating function $s$ given in 13, we have, for any $t\in[0,T]$,
@@
Integrating both sides and taking expectation in 43 yields
@@
Differentiating both sides in 44, using Young’s inequality and $m_{T-t}-m_{T-\lfloor t/\gamma\rfloor\gamma}\leq\gamma m_{T-t}$, we have
@@
We derive an upper bound for the second term on the right-hand side in 45. Using Cauchy-Schwarz inequality, Itô formula applied to $tB_{t}$ and Young’s inequality, we have
$\begin{split}&-2\mathbb{E}\left[\left\langle Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}},\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &=-2\mathbb{E}\left[\left\langle(Y_{t}^{\text{aux}}-Y_{\lfloor t/\gamma\rfloor\gamma}^{\text{aux}})-(\widehat{Y}_{t}^{\text{EM}}-\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}),\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &=-2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(-Y_{s}^{\text{aux}}+2m_{T-s}\hat{\theta})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &\quad+2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(-\widehat{Y}_{\lfloor s/\gamma\rfloor\gamma}^{\text{EM}}+2m_{T-\lfloor s/\gamma\rfloor\gamma}\hat{\theta})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &=2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(Y_{s}^{\text{aux}}-\widehat{Y}_{\lfloor s/\gamma\rfloor\gamma}^{\text{EM}})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &=2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}(Y_{s}^{\text{aux}}-Y_{\lfloor s/\gamma\rfloor\gamma}^{\text{aux}})\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &=2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(-Y_{\nu}^{\text{aux}})\mathrm{d}\nu\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &\quad+2\mathbb{E}\left[\left\langle\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\sqrt{2}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\mathrm{d}\overline{B}_{\nu}\mathrm{d}s,\sqrt{2}\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\end{split}$
$\begin{split}&\leq 2\sqrt{2}\left(\mathbb{E}\left[\left|\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(-Y_{\nu}^{\text{aux}})\mathrm{d}\nu\mathrm{d}s\right|^{2}\right]\right)^{1/2}\left(\mathbb{E}\left[\left|\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right|^{2}\right]\right)^{1/2}\\ &+4\mathbb{E}\left[\left\langle t\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}-\int_{\lfloor t/\gamma\rfloor\gamma}^{t}s\mathrm{d}\overline{B}_{s},\int_{\lfloor t/\gamma\rfloor\gamma}^{t}\mathrm{d}\overline{B}_{s}\right\rangle\right]\\ &\leq\sqrt{2}\gamma^{5/2}\left(\sup_{t\geq 0}\mathbb{E}\left[|Y_{t}^{\text{aux}}|^{2}\right]+d\right)+2d\gamma^{2}.\end{split}$ (46)
Substituting 46 into 45 yields
@@
Integrating both sides in 47 and using Lemma 16, Lemma 17 and Corollary 15, yield
@@
By using 48, we have
$W_{2}(\mathcal{L}(Y_{t_{K+1}}^{\text{aux}}),\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}))$ (49)
@@
Final upper bound on $W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}}).$
Substituting 38, 42, 49 into 34 yields
$\begin{split}&W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d})\\ &\quad+\sqrt{4/3}(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2})^{1/2}\\ &\quad+\gamma(18d+132(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2}+|\theta^{*}|^{2}))^{1/2}\\ &\leq\sqrt{2}e^{-2T}(\sqrt{\mathbb{E}\left[|X_{0}|^{2}\right]}+\sqrt{d})\\ &\quad+(\sqrt{4/3}+2\sqrt{33})(e^{-2n\lambda\mathbb{E}[\sigma_{\tau}^{2}m_{\tau}^{2}]}\mathbb{E}[|\theta_{0}-\theta^{*}|^{2}]+dC_{\mathsf{SGLD},1}/\beta+\lambda C_{\mathsf{SGLD},2})^{1/2}\\ &\quad+\gamma(18d+132|\theta^{*}|^{2})^{1/2}.\end{split}$ (50)
The bound for $W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ in 50 can be made arbitrarily small by appropriately choosing parameters including $T,\beta,\lambda,n$ and $\gamma$. More precisely, for any $\delta>0$, we first choose $T>T_{\delta}$ with $T_{\delta}$ given explicitly in Table 1 such that
@@
Using 51, 52 and 53 in 50, we obtain $W_{2}(\mathcal{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})<\delta$. ∎
C.3 Proof of the Preliminary Results
We provide the proofs of the results of Appendix C.1.
Proof of Lemma 14.
By 15, we have for any $n\in\mathbb{N}_{0}$,